The Dubs division of Fast Company (the parent company) produces wheels for off-road sport vehicles. One-half of Dub's output is sold to the Hoon division of Fast; the remainder is sold to outside customers. Dub's estimated operating profit for the year is shown in the table.

Internal Sales

External Sales

Totals

Sales

$300,000

$400,000

$700,000

Var Mfg.

$160,000

$160,000

$320,000

Var G&A

$40,000

$60,000

$100,000

CM

$100,000

$180,000

$280,000

Fixed Mfg.

$24,000

$32,000

$56,000

Fixed G&A

$36,000

$48,000

$84,000

Op. Profits

$40,000

$100,000

$140,000

Unit Sales

1,000

1,000

2,000

Unless otherwise stated assume the fixed costs given above are allocated costs and unavoidable. Hoon division has an opportunity to purchase 1,000 wheels of the same quality from an outside supplier on a continuing basis for $250.00 per wheel.

1.What is the current operating profit per unit on wheels sold internally to the Hoon division?

2.To simplify this example assume the capacity of the Dubs division is 2,000 units and it is producing and selling units as shown in the table. As a result of a marketing campaign, the Hoon division has increased its demand for wheels from Dubs to 1,400 units. The senior management of Fast company requires Dubs to provide the Hoon division with all of these units at the current internal price.
By how much would Dubs’ total contribution margin change if 1,400 units are supplied to Dubs?
Indicate an increase with a positive number and a decrease with a negative number, e.g. -10000.

3. Assume the Dubs division has excess production capacity but can only sell 1,000 units to external customers. What would be the Dubs division’s total operating profit if it agrees to meet the $250.00 price for transfers to the Hoon division?
Answer #1

1. Total Internal Unit Sales = 1,000

Total Internal Operating Profit = $40,000

Therefore, current operating profit per unit on wheels sold internally to the Hoon division = $40,000/1000 = $40

2. Total Sales for 1000 units = $300,000

Therefore, total sales for 1400 units = $(300,000/1000*1400) = $420,000

Total Var Mfg. for 1000 units = $160,000

Therefore, total var mfg. for 1400 units = $(160,000/1000*1400) = $224,000

Total Var G&A for 1000 units = $40,000

Therefore, total Var G&A for 1400 units = $(40,000/1000*1400) = $56,000

Therefore, total Contribution Margin for 1400 units = $(420,000-224,000-56,000) = $140,000

Total Contribution Margin for 1000 units = $100,000

So, Dubs Total Contribution Margin increased by $(140,000-100,000) = $40,000 if 1400 units are supplied to Hoon

3. If the Dubs division agrees to meet the $250 price for transfers to the Hoon division:

Total Sales = $250*1000units = $250,000

Var Mfg. = $160,000

Var G&A = $40,000

Fixed Mfg. = $24,000

Fixed G&A = $36,000

Therefore, Operating Profit = $(250,000-160,000-40,000-24,000-36,000) = -$10,000

So, Dubs division will have an operating loss of $10,000 if it agrees to meet the $250 price for transfers to the Hoon division